# Araida: Analogical Reasoning-Augmented Interactive Data Annotation

### v1

@@ Abstract
Chen Huang♠♣,  Yiping Jin,  Ilija Ilievski,  Wenqiang Lei♠♣thanks: Correspondence to Wenqiang Lei.,  Jiancheng Lv♠♣ College of Computer Science, Sichuan University, China NLP Group, Pompeu Fabra University, Spain Department of Industrial Systems Engineering and Management, National University of Singapore, Singapore Engineering Research Center of Machine Learning and Industry Intelligence, Ministry of Education, China huangc.scu@gmail.com, yiping.jin@upf.edu, wenqianglei@gmail.com
@@ 1 Introduction
Data annotation is a challenging task that involves a tradeoff between annotation quality and budget. While some platforms offer a cost-effective solution by relying on ML models to annotate data automatically  [^1] For example, https://aws.amazon.com/sagemaker/groundtruth/., the quality of such annotations is often compromised Wang et al. (2022a). It is particularly true in the limited data annotation scenario where the annotation budget is limited or when unlabeled data are scarce Ringger et al. (2007); Chaudhary et al. (2021).
@@ 4.1 Experimental Setup
All experiments are carried out on a machine with Intel(R) Xeon(R) Gold 5317 CPU @ 3.00GHz and a GeForce RTX 3090 GPU. For simplicity, we implement our integration strategy using a three-layer, fully-connected network with ReLu activation and dropout. For KNN, we set $k=20$ for $\rho_{t}$. KNN runs in the embedding space of text-embedding-ada-002 Neelakantan et al. (2022) when combining with ChatGPT. For Dist./FT and LLaMa2 models, KNN runs in the corresponding model’s embedding space. Moreover, we leave the details of LLM prompts and mode fine-tuning examples in Appendix B.
@@ A.3 Datastore Maintenance Strategy
Appendix B Prompts and Fine-Tuning Examples for LLM-Based Annotation Models

### v2

@@ Abstract
Chen Huang♠♣,  Yiping Jin,  Ilija Ilievski,  Wenqiang Lei♠♣thanks: Correspondence to Wenqiang Lei.,  Jiancheng Lv♠♣ College of Computer Science, Sichuan University, China Engineering Research Center of Machine Learning and Industry Intelligence, Ministry of Education, China NLP Group, Pompeu Fabra University, Spain ISEM, National University of Singapore, Singapore huangc.scu@gmail.com, yiping.jin@upf.edu, wenqianglei@gmail.com
@@ 1 Introduction
Data annotation is a challenging task that involves a tradeoff between annotation quality and budget. While some platforms offer a cost-effective solution by relying on ML models to annotate data automatically  [^1] For example, https://aws.amazon.com/sagemaker/groundtruth/., the quality of such annotations is often compromised Wang et al. (2022a). It is particularly true in the limited data annotation scenario where the annotation budget is limited or when unlabeled data are scarce Ringger et al. (2007); Chaudhary et al. (2021); Huang et al. (2024).
@@ 4.1 Experimental Setup
All experiments are carried out on a machine with Intel(R) Xeon(R) Gold 5317 CPU @ 3.00GHz and a GeForce RTX 3090 GPU. For simplicity, we implement our integration strategy using a three-layer, fully-connected network with ReLu activation and dropout. For KNN, we set $k=20$ for $\rho_{t}$. KNN runs in the embedding space of text-embedding-ada-002 Neelakantan et al. (2022) when combining with ChatGPT. For Dist./FT and LLaMa2 models, KNN runs in the corresponding model’s embedding space. Moreover, we leave the details of LLM prompts and mode fine-tuning examples in Appendix D.
@@ A.2 Top $k$ for $\rho_{t}$
Table 5: Results on comparing to fully fine-tuned models
Model WN18RR FreeBase IMDB SST-5
BERT(fully finetuned) 75.13$\pm$1.12 60.47$\pm$1.24 90.04$\pm$0.42 46.59$\pm$1.20
BERT+Araida 54.41$\pm$1.51 50.30$\pm$2.25 88.79$\pm$3.54 43.81$\pm$3.12
Dist./FT(fully finetuned) 72.64$\pm$1.05 58.50$\pm$1.36 84.87$\pm$0.81 42.12$\pm$1.44
Dist./FT+Araida 52.16$\pm$1.37 43.02$\pm$6.43 79.33$\pm$2.81 37.21$\pm$3.03
@@ A.3 Datastore Maintenance Strategy
Appendix B Results on Comparing to Fully Fine-tuned Model
We utilized up to 5K of data to simulate the limited data annotation task, and we updated the parameters of the annotation model as the interactive annotation process progressed. To validate the effectiveness of the fully fine-tuned model, we use the remaining data from the original dataset, excluding the 5K data and the validation data, as the training data to fully fine-tune the model. Next, we employ this model for interactive annotation. Since it is already fully fine-tuned, we do not update the model parameters during the annotation process. Due to time constraints, we are currently only considering scenarios where active learning has not been adopted. Additionally, we only fine-tune small annotation models (i.e., BERT and Dist./FT).
Based on the results in the Table 5, it is evident that the annotation performance of the fully fine-tuned model surpasses that of our method. This suggests that while our method currently offers a lightweight approach to aid data annotation, which can to some extent enhance the sample efficiency of the annotation model, there is still potential for improvement in future research.Nevertheless, it’s important to emphasize that in real-world data annotation scenarios, we typically don’t have access to fully fine-tuned models initially, as we lack labeled datasets. While one might turn to LLMs like ChatGPT, our findings in Figure 3 indicate that our method could potentially improve upon ChatGPT even further.
Appendix C Results on Smaller-sized Transformer-based Model
With a wide array of language models to choose from, we faced the challenge of not being able to test all available models. To address this, we selected three prominent models (GloVe, LLaMa2, and ChatGPT) based on our hardware resource capabilities in our main experiments. While open to conducting further experiments, our focus was limited by time constraints, leading us to concentrate solely on fine-tuning BERT (w/o AL), with the outcomes detailed in Table 6. After implementing Araida, we observed a significant enhancement in the quality of annotations by using our Araida.
Table 6: Results on smaller-sized Transformer-based model
Model WN18RR FreeBase IMDB SST-5
BERT 53.32$\pm$1.02 46.12$\pm$3.23 85.38$\pm$7.16 41.08$\pm$3.41
BERT+Araida 54.41$\pm$1.51 50.30$\pm$2.25 88.79$\pm$3.54 43.81$\pm$3.12
BERT+Araida w/o f 51.03$\pm$2.11 44.63$\pm$4.04 80.29$\pm$1.09 38.51$\pm$1.45
Appendix D Prompts and Fine-Tuning Examples for LLM-Based Annotation Models